OpenStudy (anonymous):

–2(6 – x) = 3x A. –3 B. 12 C. 3 D. –12

jagr2713 (jagr2713):

you distribute and get -12-(-2x)=3x then subtract -2-3=1 so -12=x

jagr2713 (jagr2713):

its D

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

jagr u still there

jagr2713 (jagr2713):

yh

OpenStudy (anonymous):

do you know the anwser

jagr2713 (jagr2713):

Its D

jagr2713 (jagr2713):

you distribute and get -12-(-2x)=3x then subtract -2-3=1 so -12=x

OpenStudy (anonymous):

Wendy earns $6.00 an hour babysitting. She babysat 4 hours Friday, 5 hours Saturday, and more hours Sunday. She earned $78.00 for the weekend. How many hours did she baby sit on Sunday? A. 4 B. 5 C. 6 D. 7

jagr2713 (jagr2713):

if he earns 6 an hour. You multiply 4*6=24 and 5*6=30 and add them and get 54. So now you just subtract 78-54=24 so 4*6=24 so A is the answer :D